OpenStudy (anonymous):

F <= (9/5)111.4 + 32 F <= 9 * 111.4/5 + 32 F <= 200.52 + 32 F <= 232.52 4 Check: C => 9/5 / 232.52 - 32 C =>9 * 232.52 / 5 - 32 C=> I need help with the check: I cant figure it out.

OpenStudy (anonymous):

Some sort of conversion for Cels -> Fahr? Boil water = 212F = 100C Freeze Water = 32F = 0C So, F = 1.8 * C + 32 and C =( F-32)/1.8

OpenStudy (anonymous):

Thanks but that's not the problem. Im trying to do the check for the f<= 232.52 Celisse is => 111.4 I got stuck on the check I'm using the 5 step method but i got the rest just need help on the check part.

OpenStudy (anonymous):

I was kinda pointing out that your equation was wrong in your cels to fahr. you used (f-32)*1.8

OpenStudy (anonymous):

o ok. i was editing it and posted I didnt mean that. i will post the problem so you can better understand

OpenStudy (anonymous):

We were given the Celsius of the temperature needed which is not more than 111.4. We must find the Fahrenheit of the temperature so that it remains stable. 2 State: Since the Celsius is less than or equal to the Fahrenheit, Using F= (9/5) C + 32 , we now have F <= (9/5)(111.4) +32

OpenStudy (anonymous):

I finally got it.......this is the answer for future references Proof F = C * 9/5 + 32 F -32 = C * 9/5 + 32 -32 - Adding the additive inverse of a number to both sides of the equation to move it to the other side of it F -32 = C * 9/5 - Addition ( F -32) * 5/9 = C * 9/5 * 5/9 - Multiplying the multiplicative inverse of a number to both sides of the equation to move it to the other side of it 5/9( F-32) = C - Multiplication C = 5/9( F- 32) - Rearranged the equation around 111.4 = 5/9 ( 232.52-32) - Substitution of C and F with 111.4 and 232.52 111.4 = 5/9( 200.52) - Subtraction 111.4 = 1,002.60 / 9 - Multiplication 111.4 = 111.4 - Division It checks and equals therefore the chemical reaction will remain stable, as long as it temperature in Farenheit is less than or equal to 232.52 degrees which is 111.4 degrees Celsius